# Graph Transformers: A Survey

### v1

@@ III-A2 Edge Structural Bias
In the realm of graph transformers, edge structural bias is crucial for extracting and understanding complex information within graph structure [52]. Edge structural bias is versatile and can represent various aspects of graph structure, including node distances, edge types, edge directions and local sub-structures. Empirical evidence has shown that edge structural encodings can improve the effectiveness of graph transformers [53, 54, 55].
Local Edge Structural Encodings. Local edge structural encodings capture the local structure of a graph by encoding relative position or distance between two nodes [52]. These encodings borrow ideas from relative positional encodings used in NLP and CV, where they are used for modeling sequential or spatial order of tokens or pixels [56]. However, in the context of graphs, the concept of relative position or distance between nodes becomes ambiguous due to the presence of multiple connecting paths with varying lengths or weights. Consequently, the scientific community has proposed various methods to define and encode this information specifically for graph structures.
GraphiT [46] introduces local edge structural encodings to graph transformers. It promotes the use of positive definite kernels on graphs to measure node similarity considering the shortest path distance between them. The kernel function is defined as:
$k(u,v)=\exp(-\alpha d(u,v)),$ (12)
where $u$ and $v$ are two nodes in a graph, $d(u,v)$ is their shortest path distance and $\alpha$ is a hyperparameter that controls decay rate. The kernel function is then used to modify the self-attention score between two nodes as:
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+k(\textbf{Q},\textbf{K}))\textbf{V},$ (13)
where $k(\textbf{Q},\textbf{K})$ is the matrix of kernel values computed for each pair of nodes. EdgeBERT [57] proposes using edge features as additional input tokens for graph transformers. These edge features are obtained by applying a learnable function to the source and target node features of each edge. The resulting edge features are then concatenated with node features and fed into a standard transformer encoder.
More recently, the Edge-augmented Graph Transformer (EGT) [58] introduces residual edge channels as a mechanism to directly process and output both structural and node information. The residual edge channels are matrices that store edge information for each pair of nodes. They are initialized with either an adjacency matrix or the shortest path matrix and updated at each transformer layer by applying residual connections. These channels are then used to adjust the self-attention score between two nodes
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V},\textbf{R}_{e})=\text{softmax}(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+\textbf{R}_{e})\textbf{V},$ (14)
where $\textbf{R}_{e}$ is the residual edge channel matrix.
Although local edge structural encodings can capture detailed structural information, they may have limitations in capturing overall structural information. This can lead to increased computational complexity or memory usage as pairwise information needs to be computed and stored. Additionally, the effectiveness of these encodings may vary depending on the selection and optimization of encoding or kernel function for different graphs and tasks.
Global Edge Structural Encodings. Global edge structural encodings aim to capture the overall structure of a graph. Unlike NLP and CV domains, the exact position of a node in graphs is not well-defined because there is no natural order or coordinate system [59]. Several approaches have been suggested to tackle this issue.
@@
GPT-GNN [60] is an early work that utilizes graph pooling and unpooling operations to encode the hierarchical structure of a graph. It reduces graph size by grouping similar nodes and then restoring the original size by assigning cluster features to individual nodes. This approach results in a multiscale representation of graphs and has demonstrated enhanced performance on diverse tasks. Graphormer [11] uses spectral graph theory to encode global structure. It uses eigenvectors of normalized Laplacian matrix as global positional encodings for nodes. This method can capture global spectral features (e.g., connectivity, centrality and community structure). Park et al. [41] extended Graphormer by using singular value decomposition (SVD) to encode global structure. They utilized the left singular matrix of the adjacency matrix as global positional encodings for nodes. This approach can handle both symmetric and asymmetric matrices.
Global edge structural encodings excel at capturing coarse-grained structural information at the graph level, benefiting tasks that require global understanding. However, they may struggle with capturing fine-grained node-level information and can lose data or introduce noise during encoding. In addition, their effectiveness may depend on the choice of encoding technique and matrix representations.
@@ III-A3 Message-passing Bias
Message-passing bias is a crucial inductive bias for graph transformers to facilitate learning from the local structure of graphs [61]. This bias enables graph transformers to exchange information between nodes and edges, thereby capturing dependencies and interactions among graph elements. Moreover, message-passing bias helps graph transformers in overcoming certain limitations of standard transformer architecture, such as the quadratic complexity of self-attention, the absence of positional information and challenges associated with handling sparse graphs [39]. Formally, message-passing bias can be expressed as follows:
$\textbf{h}_{v}^{(t+1)}=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$ (15)
Here, $\textbf{h}_{v}^{(t)}$ is the representation of node $v$ at layer $t$, $u$ is a neighboring node of node $v$, $\textbf{e}_{uv}$ is the feature of the edge between nodes $v$ and $u$, and $f$ is the function that aggregates information from neighbors and edges. Incorporation of message-passing bias in graph transformers can be achieved through three primary approaches: preprocessing, interleaving and post-processing (as illustrated in Figure 4 ). These approaches vary in how they combine message-passing operations with self-attention layers within transformer architecture.
@@
Preprocessing. Preprocessing involves applying message-passing operations to node features before feeding them to self-attention layers [36]. This technique aims to augment node features with local structural information, making them more compatible with global self-attention. Preprocessing can maintain the integrity of the original transformer architecture without making any modifications and utilize message-passing modules from existing GNNs. It can be mathematically defined as follows:
@@
$\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}).$
Here, $\textbf{h}_{v}^{(0)}=\textbf{x}_{v}$, $\textbf{x}_{v}$ is input feature of node $v$, SelfAttention is a function that performs self-attention over all nodes.
The limitation of preprocessing is that it applies message-passing only once before self-attention layers. This approach may not fully capture intricate interactions between nodes and edges on various scales. Additionally, the preprocessing step may introduce redundancy and inconsistency between the message-passing module and the self-attention layer as they both serve similar functions of aggregating information from neighboring elements.
Interleaving. Interleaving refers to the technique employed in graph transformer architecture that involves alternating message-passing operations and self-attention layers [62, 16, 63]. The objective of this technique is to achieve a balance between local and global information processing, thereby enabling multi-hop reasoning over graphs. By integrating message-passing modules into core components of graph transformers, interleaving enhances their expressive power and flexibility. It can be mathematically defined as follows:
@@
$\theta$ $=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$
One drawback of the interleaving technique is its impact on the complexity and computational requirements of graph transformers. This is due to the need for additional parameters and operations compared to pre-processing or post-processing. Furthermore, interleaving can potentially lead to interference and conflict between the message-passing module and the self-attention layer as they each update node representations in distinct manners.
Post-processing. Post-processing refers to the technique of applying message-passing operations to node representations obtained from self-attention layers [64, 58]. The purpose of this approach is to refine and adjust node representations based on underlying graph structure, thereby enhancing their interpretability and robustness. By doing so, this method aims to improve the quality and utility of node representations for downstream tasks and applications. It can be mathematically defined as follows:
@@
$\textbf{h}_{v}^{(T+1)}$ $=f(\textbf{h}_{v}^{(T)},\{\textbf{h}_{u}^{(T)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}).$
Here, $T$ is the final layer of the graph transformer.
The drawback of post-processing is its limited application of message-passing after self-attention layers, potentially failing to capture intricate semantics and dynamics of graph data. Additionally, post-processing runs the risk of introducing noise and distortion to node representations as it has the potential to overwrite or override information acquired by self-attention layers.
@@ III-A4 Attention Bias
Attention bias enables graph transformers to effectively incorporate graph structure information into the attention mechanism without message-passing or positional encodings [39, 65]. Attention bias modifies attention scores between nodes based on their relative positions or distances in the graph. It can be categorized as either local or global depending on whether it focuses on the local neighborhood or global topology of the graph.
Local Attention Bias. Local attention bias limits the attention to a local neighborhood surrounding each node. This concept is analogous to the message-passing mechanism observed in GNNs [66]. It can be mathematically defined as follows:
[…]
where $\alpha_{ij}$ is the attention score between node $v_{i}$ and node $v_{j}$, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are their node features, $g$ is a function that computes the similarity between two nodes, such as dot-product and linear transformation. $b_{ij}$ is a local attention bias term that modifies attention score based on the distance between node $v_{i}$ and node $v_{j}$. The local attention bias term can be either a binary mask that only allows attention within a certain hop distance [67, 42, 36] or a decay function that decreases attention score with increasing distance [68, 64].
Global Attention Bias. Global attention bias integrates global topology information into the attention mechanism independent of message-passing and positional encodings [69]. It can be mathematically defined as follows:
@@ III-B Graph Attention Mechanisms
Graph attention mechanisms play an important role in the construction of graph transformers [71]. By dynamically assigning varying weights to nodes and edges in the graph, these mechanisms enable transformers to prioritize and emphasize the most relevant and important elements for a given task [72]. Specifically, a graph attention mechanism is a function that maps each node $v_{i}\in V$ to a vector $\textbf{h}_{i}\in\mathbb{R}^{d_{k}}$:
$\textbf{h}_{i}=f_{n}(\textbf{x}_{i},\{\textbf{x}_{j}\}_{v_{j}\in\mathcal{N}(v_{i})},\{A_{ij}\}_{v_{j}\in\mathcal{N}(v_{i})}),$ (21)
where $f_{n}$ is a nonlinear transformation, $\textbf{x}_{i}$ is the input feature vector of node $v_{i}$, $\textbf{x}_{j}$ is the input feature vector of node $v_{j}$. The function $f_{n}$ can be decomposed into two parts: an attention function and an aggregation function.
The attention function computes a scalar weight for each neighbor of node $v_{i}$, denoted by $\alpha_{ij}$, which reflects the importance or relevance of node $v_{j}$ for node $v_{i}$:
$\alpha_{ij}=\text{softmax}_{i}(\text{LeakyReLU}(\textbf{W}_{a}[\textbf{x}_{i}||\textbf{x}_{j}])),$ (22)
where $\textbf{W}_{a}\in\mathbb{R}^{1\times 2d_{n}}$ is a learnable weight matrix, ||||| | is the concatenation operation, and $\text{softmax}_{i}$normalizes weights over all neighbors of node $v_{i}$. The aggregation function combines the weighted features of the neighbors to obtain the output representation of node $v_{i}$:
@@
where $\textbf{W}_{h}$ is another learnable weight matrix. Graph attention mechanisms have the potential to extend their application beyond nodes to edges. This can be achieved by substituting nodes with edges and utilizing edge features instead of node features [73]. Furthermore, stacking of graph attention mechanisms in multiple layers allows for the utilization of output representations from one layer as input features for the subsequent layer [74].
@@ III-B1 Global Attention Mechanisms
Global attention mechanisms can determine how each node calculates its attention weights across all other nodes in the graph [75]. Global attention mechanisms can be broadly categorized into two types: quadratic attention mechanisms and linear attention mechanisms.
Quadratic Attention Mechanisms. Quadratic attention mechanisms are derived from the conventional self-attention formula. This formula calculates attention weights by applying a softmax function to scale the dot product between the query and key vectors of each node:
$\alpha_{ij}=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}})}{\sum_{n=1}^{N}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}})}.$ (24)
The computational complexity of this method is $O(N^{2})$.
One of the pioneering studies that has introduced quadratic attention mechanisms for graph transformers is the work by Velickovic et al.[42]. The authors proposed the use of multiple attention heads to capture different types of relations between nodes. Choromanski et al.[76] introduced the Graph Kernel Attention Transformer (GKAT), an approach that integrates graph kernels, structural priors, and efficient transformer architectures. GKAT leverages graph kernels as positional encodings to capture the structural characteristics of the graph while employing low-rank decomposition techniques to minimize the memory requirements of quadratic attention:
$\alpha_{ij}$ $=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{j})}{\sum_{n=1}^{N}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{n})},$ (25)
$\mathbf{z}_{i}$ $=\sum_{j=1}^{N}\alpha_{ij}(\mathbf{U}\mathbf{V}^{\top})_{ij}$
where $\textbf{z}_{i}$ is the output vector of node $v_{i}$, $\mathbf{p}_{i}$ is the positional encoding vector of node $v_{i}$, which is computed by applying a graph kernel function to node features. $\mathbf{U}$ and $\mathbf{V}$ are low-rank matrices that approximate value matrix in the GKAT model. GKAT also introduced a novel kernel-based masking scheme to control the sparsity of the attention matrix. Yun et al. [36] introduced Graph Transformer Networks (GTN), an approach that uses multiple layers of quadratic attention to acquire hierarchical representations of graphs. GTN further enhances attention computation by incorporating edge features through the use of bilinear transformations instead of concatenation. This innovative technique demonstrates the potential for improved graph representation learning in the field of graph analysis.
Quadratic attention mechanisms possess a significant advantage in capturing extensive dependencies and global information within graphs, thereby offering potential benefits for various graph learning tasks. Nonetheless, these mechanisms have some limitations, including their computationally expensive nature, high memory consumption, and susceptibility to noise and outliers. Additionally, quadratic attention mechanisms may not effectively preserve and leverage local information within graphs which holds significance in certain domains and tasks.
Linear Attention Mechanisms. Linear attention mechanisms employ different approximations and hashing techniques to decrease the computational complexity of self-attention from $O(N^{2})$ to $O(N)$ [77]. These mechanisms can be categorized into two subtypes: kernel-based linear attention mechanisms and locality-sensitive linear attention mechanisms.
@@
Kernel-based linear attention mechanisms leverage the concept of employing kernel functions to map query and key vectors into a feature space that allows efficient computation of their inner products [78, 79]. In the work by Katharopoulos et al. [80], Linear Transformer Networks (LTN) were introduced. LTN utilizes random Fourier features as kernel functions to approximate self-attention:
$\alpha_{ij}$ $=\frac{\exp(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{j})}{\sqrt{d_{k}}})}{\sum_{n=1}^{N}\exp(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{n})}{\sqrt{d_{k}}})},$ (26)
$\phi(\mathbf{x})$ $=\sqrt{\frac{2}{m}}[\cos(\omega_{1}^{\top}\mathbf{x}+b_{1}),...,\cos(\omega_{m}^{\top}\mathbf{x}+b_{m})]^{\top}.$
Here, $\phi$ is a random Fourier feature function that maps input vector into a $m$-dimensional feature space. $\omega_{i}$ and $b_{i}$ are randomly sampled from a Gaussian distribution and a uniform distribution, respectively.
Locality-sensitive linear attention mechanisms are based on the concept of employing hashing functions to partition the query and key vectors into discrete buckets that facilitate local computation of their inner products [81, 82, 83]. Kitaev et al. [84] introduced the Reformer model which uses locality-sensitive hashing (LSH) as a hashing function to approximate self-attention:
$\alpha_{ij}$ $=\frac{\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}})}{\sum_{n\in B_{i}}\exp(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}})},$ (27)
$B_{i}$ =$=\{j|h(\mathbf{q}_{i})=h(\mathbf{k}_{j})\},$
where $h$ is an LSH function that maps the input vector into a bucket index and $B_{i}$ is a set of nodes that share the same bucket index with node $v_{i}$. RGN also incorporated edge features into the attention computation using a shared query-key projection matrix.
Linear attention mechanisms offer a significant advantage in terms of reducing computational cost and memory usage in self-attention, thereby enabling graph transformers to efficiently process large-scale graphs. However, they do come with certain limitations, including approximation error, hashing collision and loss of global information. Additionally, linear attention mechanisms may not be capable of capturing intricate or nonlinear relationships between nodes.

### v2

@@ III-A2 Edge Structural Bias
In graph transformers, edge structural bias provides a way to extract and interpret complex graph structure information [47]. It can capture diverse structural aspects, including node distances, edge types, edge directions, and local sub-structures. Studies indicate that edge structural encoding can improve the performance of graph transformers [48, 49, 50].
Local Edge Structural encoding. Local edge structural encoding captures the local structure of a graph by encoding the relative position or distance between nodes [47]. Drawing on relative positional encoding in NLP and CV [51], these methods must account for the fact that graph distances can be ambiguous due to multiple paths with varying lengths or weights. GraphiT [41] introduces local edge structural encoding by using positive definite kernels to measure node similarity based on shortest path distance: $k(u,v)=\exp(-\alpha d(u,v)),$ where $u$ and $v$ are nodes, $d(u,v)$ is the shortest path distance, and $\alpha$ controls the decay rate. This kernel modifies the self-attention score: $\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+k(\textbf{Q},\textbf{K})\right)\textbf{V},$ where $k(\textbf{Q},\textbf{K})$ is the matrix of kernel values. EdgeBERT [52] uses edge features as additional input tokens, obtained by applying a learnable function to source and target node features of each edge, then concatenating them with node features for transformer processing. The Edge-augmented Graph Transformer (EGT) [53] introduces residual edge channels, matrices that store edge information for each node pair. Initialized with an adjacency matrix or shortest path matrix, these channels are updated at each transformer layer using residual connections to adjust self-attention scores: $\text{Attention}(\textbf{Q},\textbf{K},\textbf{V},\textbf{R}_{e})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}+\textbf{R}_{e}\right)\textbf{V},$ where $\textbf{R}_{e}$ is the residual edge channel matrix. Local edge structural encoding can capture fine-grained structural information but may overlook global graph properties, and it may increase computational cost depending on the chosen kernel or encoding strategy.
@@
Global Edge Structural encoding. Global edge structural encoding aims to represent the overall structure of a graph, which is difficult due to the absence of a natural order or coordinate system [54]. GPT-GNN [55] uses graph pooling and unpooling operations to encode hierarchical structures, reducing the graph size by grouping similar nodes and then restoring the original size by assigning cluster features back to individual nodes. Graphormer [11] applies spectral graph theory, using eigenvectors of the normalized Laplacian matrix as global positional encoding to capture global spectral features such as connectivity and community structure. Park et al. [36] extended this by using singular value decomposition (SVD), employing the left singular matrix of the adjacency matrix as global positional encoding, suitable for both symmetric and asymmetric matrices. Global edge structural encoding effectively captures coarse-grained structure at the graph level but may lose detail about individual nodes or local patterns, and its success depends on the chosen encoding technique and matrix representations.
@@ III-A3 Message-passing Bias
Message-passing bias enables graph transformers to utilize local graph structure by exchanging information between nodes and edges, addressing certain limitations of the standard transformer architecture, such as the quadratic complexity of self-attention and the lack of positional information [56, 34]. Approaches to incorporate this bias in graph transformers include preprocessing, interleaving, and post-processing, each employing distinct methods of combining message-passing operations with self-attention layers (see Figure 3).
@@
Preprocessing. Preprocessing applies message-passing operations to node features before self-attention layers [31], thereby augmenting node features with local structural information. This approach retains the original transformer design and employs standard GNN-based message-passing modules. Mathematically, it is defined as:
@@
$\textbf{h}_{v}^{(t+1)}$ $=\text{SelfAttention}(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in V\}),$
where $\textbf{h}_{v}^{(0)}=\textbf{x}_{v}$. Although this method provides local structural cues, applying message-passing only once before self-attention layers may limit capturing multi-scale interactions. It can also introduce redundancy between the message-passing module and the self-attention layer.
Interleaving. Interleaving alternates message-passing operations and self-attention layers [57, 16, 58], balancing local and global information processing. This arrangement supports multi-hop reasoning and can improve the expressive power of graph transformers. Mathematically, it is expressed as:
@@
$\theta$ $=f(\textbf{h}_{v}^{(t)},\{\textbf{h}_{u}^{(t)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}),$
where added parameters and operations can increase computational cost, and conflicting updates by message-passing and self-attention may require careful tuning.
Post-processing. Post-processing applies message-passing to the node representations produced by self-attention layers [59, 53], refining them based on the underlying graph structure. Mathematically:
@@
$\textbf{h}_{v}^{(T+1)}$ $=f(\textbf{h}_{v}^{(T)},\{\textbf{h}_{u}^{(T)}:u\in\mathcal{N}(v)\},\{\textbf{e}_{uv}:u\in\mathcal{N}(v)\}),$
where $T$ denotes the final layer of the graph transformer. Since message-passing occurs only after self-attention, it may overlook complex interactions at earlier layers. Furthermore, overwriting information from self-attention layers can introduce inconsistencies in the final node representations.
@@ III-A4 Attention Bias
Attention bias allows graph transformers to incorporate graph structure information into the attention mechanism without needing message-passing or positional encoding [34, 60]. It modifies attention scores between nodes based on their relative positions or distances in the graph and can be categorized as either local or global, depending on whether it focuses on the local neighborhood or global topology of the graph.
Local Attention Bias. Local attention bias restricts attention to a node’s local neighborhood, similar to the message-passing mechanism in GNNs [61]. It is mathematically defined as:
$\alpha_{ij}=\frac{\exp(g(\mathbf{x}_{i},\mathbf{x}_{j})\cdot b_{ij})}{\sum_{k\in\mathcal{N}(v_{i})}\exp(g(\mathbf{x}_{i},\mathbf{x}_{k})\cdot b_{ik})},$ (6)
where $\alpha_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are their node features, and $g$ computes node similarity, such as dot-product or linear transformation. The term $b_{ij}$ is a local attention bias that adjusts the attention score based on the distance between nodes $v_{i}$ and $v_{j}$. This bias can be a binary mask allowing attention within a certain hop distance [62, 37, 31] or a decay function decreasing attention with increasing distance [63, 59].
Global Attention Bias. Global attention bias integrates global topology information into the attention mechanism independent of message-passing and positional encoding [64]. It can be mathematically defined as follows:
@@ III-B Graph Attention Mechanisms
Graph attention mechanisms are widely used in constructing graph transformers [66]. By dynamically weighting nodes and edges, they enable transformers to focus on the most relevant parts of a graph for a given task [67]. Formally, a graph attention mechanism maps each node $v_{i}\in V$ to a vector $\textbf{h}_{i}\in\mathbb{R}^{d_{k}}$: $\textbf{h}_{i}=f_{n}(\textbf{x}_{i},\{\textbf{x}_{j}\}_{v_{j}\in\mathcal{N}(v_{i})},\{A_{ij}\}_{v_{j}\in\mathcal{N}(v_{i})}),$ where $f_{n}$ is a nonlinear function, $\textbf{x}_{i}$ is the feature of node $v_{i}$, and $\textbf{x}_{j}$ is the feature of neighbor $v_{j}$. The mechanism typically comprises an attention function and an aggregation function. The attention function computes a scalar weight $\alpha_{ij}$ for each neighbor: $\alpha_{ij}=\text{softmax}_{i}(\text{LeakyReLU}(\textbf{W}_{a}[\textbf{x}_{i}||\textbf{x}_{j}])),$ where $\textbf{W}_{a}\in\mathbb{R}^{1\times 2d_{n}}$ is a learnable parameter and $||$ denotes concatenation. The aggregation function then combines weighted neighbor features:
@@
where $\textbf{W}_{h}$ is another learnable matrix. These mechanisms can also be applied to edges by treating edges as primary elements and using edge features [68]. Stacking multiple layers of graph attention allows each layer’s output to inform subsequent layers [69].
@@ III-B1 Global Attention Mechanisms
Global attention mechanisms define how each node allocates attention weights to all other nodes in the graph [70]. They are commonly divided into two main categories: quadratic attention and linear attention.
Quadratic Attention. Quadratic attention follows the standard self-attention formulation, using a softmax function on the scaled dot product of query and key vectors:
$\alpha_{ij}=\frac{\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}\bigr)}{\sum_{n=1}^{N}\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}\bigr)},$ (9)
with a computational complexity of $O(N^{2})$. Veličković et al. introduced this approach for graph transformers with multi-head attention [37]. Choromanski et al. proposed the Graph Kernel Attention Transformer (GKAT), which combines graph kernels, structural priors, and efficient attention:
$\alpha_{ij}$ $=\frac{\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{j}\bigr)}{\sum_{n=1}^{N}\exp\bigl(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}+\mathbf{p}_{i}^{\top}\mathbf{p}_{n}\bigr)},$ (10)
$\mathbf{z}_{i}$ $=\sum_{j=1}^{N}\alpha_{ij}(\mathbf{U}\mathbf{V}^{\top})_{ij},$
where $\mathbf{p}_{i}$ is a positional encoding based on graph kernels, and $\mathbf{U},\mathbf{V}$ are low-rank matrices approximating the value matrix [71]. Yun et al. introduced Graph Transformer Networks (GTN), which compute hierarchical representations using multiple layers of quadratic attention [31]. Despite their ability to capture global relationships, these mechanisms can be memory-intensive and prone to noise, and they may overlook some local information.
@@
Linear Attention. Linear attention mechanisms lower self-attention’s complexity from $O(N^{2})$ to $O(N)$ by using approximations or hashing [72]. They typically fall into two subtypes: kernel-based and locality-sensitive. Kernel-based methods map query-key vectors into a feature space for efficient dot-product computation [73, 74]. For instance, Katharopoulos et al. introduced Linear Transformer Networks (LTN) with random Fourier features [75]:
$\alpha_{ij}$ $=\frac{\exp\bigl(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{j})}{\sqrt{d_{k}}}\bigr)}{\sum_{n=1}^{N}\exp\bigl(\frac{\phi(\mathbf{q}_{i})^{\top}\phi(\mathbf{k}_{n})}{\sqrt{d_{k}}}\bigr)},$ (11)
$\phi(\mathbf{x})$ $=\sqrt{\frac{2}{m}}\bigl[\cos(\omega_{1}^{\top}\mathbf{x}+b_{1}),\dots,\cos(\omega_{m}^{\top}\mathbf{x}+b_{m})\bigr]^{\top}.$
Locality-sensitive linear attention uses hashing to group queries and keys into buckets for local dot-product computation [76, 77, 78]. Kitaev et al.’s Reformer model approximates self-attention with locality-sensitive hashing [79]:
$\alpha_{ij}$ $=\frac{\exp\bigl(\tfrac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d_{k}}}\bigr)}{\sum_{n\in B_{i}}\exp\bigl(\tfrac{\mathbf{q}_{i}^{\top}\mathbf{k}_{n}}{\sqrt{d_{k}}}\bigr)},$ (12)
$B_{i}$ $=\{j\mid h(\mathbf{q}_{i})=h(\mathbf{k}_{j})\},$
where $h$ is a locality-sensitive hash function. Linear attention mechanisms are more scalable for large graphs but can introduce approximation errors, hashing collisions, and partial loss of global information, potentially limiting their ability to capture complex node relationships.